According to Masterfoods, the company that manufactures M&M’s, 12% of peanut M&M’s are brown, 15% are yellow, 12% are red, 23% are blue, 23% are orange and 15% are green. You randomly select peanut M&M’s from an extra-large bag looking for a blue candy. (Round all probabilities below to four decimal places; i.e. your answer should look like 0.1234, not 0.1234444 or 12.34%.)

Compute the probability that the first blue candy is the eighth M&M selected.



Compute the probability that the first blue candy is the eighth or ninth M&M selected.



Compute the probability that the first blue candy is among the first eight M&M’s selected.



If every student in a large Statistics class selects peanut M&M’s at random until they get a blue candy, on average how many M&M’s will the students need to select? (Round your answer to two decimal places.)
Answer #1
this is Geometric distribution with parameter p=0.23

a)

probability that the first blue candy is the eighth M&M selected :

P(X=8)= (1-p)x-1p= 0.0369

b)

probability that the first blue candy is the eighth or ninth M&M selected :

P(8<=X<=9)= (1-p)x1-1-(1-p)x2= 0.0653

c)

probability that the first blue candy is among the first eight M&M’s selected :

P(X<=8)= 1-(1-p)x= 0.8764

d)

expected number =1/p=1/0.23 =4.35